OpenStudy (anonymous):

Carla has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 10 inches long, and 7 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? A) The paint will not fill the tray by 44.50 in3. B)The paint will not fill the tray by 469 cm3 C)The paint will fill the paint exactly D)The paint will overfill the tray by 44.59 in3.

OpenStudy (mathstudent55):

Let's find the volume of the tray in in^3. \(V = LWH = 10~in. \times 10~in. \times 7~cm \times \dfrac{1~in.}{2.54~cm} = 275.59~in.^3\) Now you have the volume of the tray in in.^3. Since the volume of the tray (275.59 in.^3) is greater than the volume of the paint (1 gallon = 23 in.^3), the paint will not fill the tray. The amount of paint missing is the volume of the tray minus the volume of the paint.

OpenStudy (anonymous):

Thanks